Engineering economics
aste site for a city must be located at Site A or Site B. After sorting e of the solid refuse will be transported to an electric power plant where it will be Q2. A municipal solid-w som used as fuel. Data for the hauling of refuse from each site to the power plant are shown in Table below: TABLE P2-4 Table for Problem 2-4 Site A Site B Average hauling distance 4 miles 3 miles ︼-Annual rental fee for solid-waste site Hauling cost $10,00 $90,00 $1.50/yd3-mile $1.50/yd3-mile If the power plant will pay $7.50 per cubic yard of sorted solid waste delivered to the plant, a) At which site he solid-waste should be located? Use the citys viewpoint and assume that 190,000 cubic yards of refuse will be hauled to the plant for one year only. Justify your answer. (16 Marks) b) For the selected site, what is the breakeven demand in terms of delivered cubic yard of paving materials? (8 Marks)
Answer #1

a)Total cost of hauling of refuse from site A to power plant = Hauling cost/yd3 -mile * total solid waste *distance between Site A annd power plant + Annual Rental fee for solid waste site = 1.50*190000*4 +1000= $11,41,000

Similarly, total cost of hauling of solid waste refuse from site B to power plant = 1.50*190000*3 +9000= $86,4000

Also total revenue from power plant for hauling of solid waste = 7.5*190000= $14,25,000 As total revenue of the city for hauling 90,000 cubic yards of solid waste from site A or site B to power plant is same.

By comparing total hauling cost of solid wastes for site A and site B, we find overhauling cost of solid waste for site B is lower than Site A. Therefore solid wastes should be located in Site B.

b) Breakeven demand is the demand at breakeven point where total revenue-total cost = 0

Here total cost for the selected firm is $86,4000. therefore total revenue at breakeven point is also = $86,4000

Also Average revenue = $7.50/per cubic yard of solid waste

Therefore breakeven demand of delivered solid waste refuse for the selected site = 864000/7.50 = 11,5200 cubic yards.